OpenStudy (anonymous):

solve for the variable 20^(-6n)+6=55 please check my work

OpenStudy (ageta):

20−6n+6=55 Step 1: Add -6 to both sides. 20−6n+6+−6=55+−6 20−6n=49 Step 2: Solve Exponent. 20−6n=49 log(20−6n)=log(49)(Take log of both sides)

OpenStudy (ageta):

n=−0.21652

OpenStudy (anonymous):

20^(-6n)=49 -6nlog(20)=log(49) -6n=log(49)/log(20) n=